# Theoretical Modeling of LLM Self-Improvement Training Dynamics Through Solver-Verifier Gap

### v3

@@ Abstract
Self-improvement is among the most prominent techniques within the realm of large language model (LLM), aiming to enhance the LLM performance without relying on external data. Despite its significance, generally how LLM performance evolves during the self-improvement process remains underexplored. In this paper, we theoretically model the training dynamics of self-improvement via the concept of solver-verifier gap. This is inspired by the conjecture that the performance enhancement of self-improvement stems from the gap between LLM’s solver capability and verifier capability. Based on the theoretical framework, we further show how to model the entire training trajectory. This framework allows quantifying the capability limit of self-improvement by fitting the theoretical model to the experiment results. We validate the effectiveness of the theoretical framework on various LLMs and datasets. Beyond self-improvement, we extend our analysis to investigate how external data influences these dynamics within the framework. Notably, we find that under limited external data regimes, such external data can be utilized at any stage without significantly affecting final performances, which accords with the empirical observations.
{NoHyper}**footnotetext: Equal contributions.
@@ 1 Introduction
In this paper, we theoretically model the training dynamics of LLM self-improvement, inspired by the conjecture that self-improvement capability arises from the gap between an LLM’s solver capability and verifier capability (Song et al., 2025). Specifically, we define these two capabilities as follows:
@@
where $\alpha^{\prime},\beta^{\prime},k$ denotes the constants with detailed formulations in Section 3.2, and $U_{s,\infty},U_{v,\infty}$ represent the solver capability and the verifier capability at convergence, respectively. Along the trajectory, both capabilities follow exponential convergence laws according to the framework. Besides, we specifically focus on the solver’s ultimate capability $U_{s,\infty}=\frac{1}{\alpha-\beta}(\alpha U_{v,0}-\beta U_{s,0}+\alpha\frac{b}{k})$. This result reveals that the solver’s ultimate capability relates to both the initial capability $U_{s,0},U_{v,0}$ and the coefficient $\alpha,\beta$. Therefore, the ultimate capability might be improved given a larger verifier-solver gap at initialization, in accordance with our insights. Detailed derivations are provided in Section 3.2.
@@
From the experimental perspective, we find in Section 4 that such theoretical modeling demonstrates strong practical efficacy. Specifically, experimental results across various models and datasets in Figure 3 reveal that the capability dynamics indeed follow an exponential law, as implied by the theoretical framework in Section 3. Besides, we observe in Figure 3 that the verifier capability consistently outperforms the solver throughout the self-improvement process, which might be empirical evidence of the solver-verifier gap’s crucial role in driving capability improvement. To further investigate the role of the solver-verifier gap, we conduct experiments in Section 4.3 on multiple LLMs demonstrating that the gap generally occurs in practice. These findings hold across varying sample sizes and even under cross-evaluation scenarios.
@@
We further analyze in Section 5 the application of this framework in cross-improvement, a potential approach to enhance the ultimate capability of self-improvement. Cross-improvement in this paper refers to the utilization of external data in the verification step (details in Figure 6). Within the theoretical framework, we contend that cross-improvement outperforms self-improvement due to the enhanced verification capabilities. We then derive the training dynamics under the cross-improvement regimes, with enhanced verification capabilities compared to self-improvement. This analysis can further assist in answering the question: given the limited amount of external data, how should one allocate these external data during the cross-improvement training process? Our theoretical analyses demonstrate that the timing of using external data is not crucial; rather, the utilization of such data genuinely enhances the training process. Therefore, external data can be incorporated at any stage as desired. Experiments in Section 5.2 on the cross-improvement validate the theoretical findings. Our main contributions are
  • Theoretical framework: We detail our theoretical framework on the dynamics of self-improvement in Section 3, inspired by the potential energy framework based on the solver-verifier capability gap in Equation 7. Theoretical derivations in Equation 12 imply an exponential law for the model capability, and we further establish the extreme capability based on the framework.
  • Experiments: We conduct experiments on self-improvement to verify the theoretical framework in Section 4 across multiple models, datasets, and verification methods. Our empirical observations indicate that (i) the uncertainty/accuracy during the self-improvement process follows an exponential law with an extreme capability (Figure 3) as implied by the theoretical framework; (ii) the solver-verifier gap generally happens in practice across different regimes (Section 4.3), validating the utility of the gap as the potential energy.
  • Cross-improvement: We further investigate in Section 5 the cross-improvement under the above framework, where limited external data is provided during the training process. We contend in Section 5.1 that cross-improvement might improve the verifier capability, thus improving the extreme capability of self-improvement. Empirically, we observe in Section 5.2 that if the verifier capability is improved by the external data, the performances of LLM are enhanced, and vice versa (Table 1). This accords with the theoretical findings.
@@ 2 Related Work
Self-Improvement. Self-improvement aims to enhance model performance without relying on external information (Huang et al., 2023; Wang et al., 2023; Bai et al., 2022; Pang et al., 2024). Self-improvement is of significant importance as it enables models to adapt and evolve autonomously, thereby facilitating their effectiveness across a wide range of real-world scenarios, such as reasoning (Zelikman et al., 2022; Peng et al., 2024; Huang et al., 2023), alignment (Wang et al., 2023, 2024b; Ding et al., 2024), and planning (Tian et al., 2024). In this paper, we consider a branch of self-improvement that fine-tunes with the output of LLM (Amini et al., 2024; Sessa et al., 2024; Gui et al., 2024; Pace et al., 2024; Ouyang et al., 2022; Rafailov et al., 2023). Alternative approaches to self-improvement include self-distillation (Buciluǎ et al., 2006; Hinton et al., 2015; Zhang et al., 2019) which involves transferring knowledge from a larger, more complex model to a smaller one, and self-correction (Kumar et al., 2024; Liu et al., 2024) where the model identifies and rectifies its own errors, etc. A body of research has also explored the potential negative consequences of self-improvement, including degradation issues (Bertrand et al., 2024; Gerstgrasser et al., 2024) and failures on out-of-domain reasoning tasks (Yuan et al., 2025).
Theoretical Understandings on Self-Improvement. Theoretical insights into self-improvement could potentially enhance comprehension, thereby making self-improvement more reliable (Yampolskiy, 2015). Previous works have theoretically studied self-improvement via self-distillation techniques, providing convergence rates for linear models (Mobahi et al., 2020; Frei et al., 2022; Das and Sanghavi, 2023; Pareek et al., 2024), neural networks (Allen-Zhu and Li, 2023), and general models (Boix-Adsera, 2024). In the realm of LLM, several works have theoretically explored self-improvement with in-context alignment (Wang et al., 2024c), reinforcement learning (Talvitie, 2017; Choi et al., 2024; Gandhi et al., 2025), meta learning (Kirsch and Schmidhuber, 2022), and diffusion models (Fu et al., 2024). Nevertheless, the theoretical understanding of self-improvement in the context of LLM training dynamics remains underexplored.
@@ 3 Theoretical Modeling of Self-Improvement Training Dynamics
This section presents a theoretical framework for sketching training dynamics in LLM self-improvement. We start by introducing necessary notations of self-improvement in Section 3.1. We then theoretically model the self-improvement dynamics in Section 3.2.
@@ 3.1 Solver and Verifier
Notations. Let $(x,y)$ denote a prompt-response pair, with $y^{[k]}$ denoting the $k$-th token and $y^{[1:k]}$ denoting the first $k$ tokens. We use $L(y)$ to represent the length of $y$. Let $\pi_{f}(y|x)$ denote the probability that a model $f$ generates a response $y$ given a prompt $x$, where $\pi_{f}(y|x)$ can be split with the auto-regressive structure of the response $\pi_{f}(y|x)=\prod_{k=1}^{L(y)}\pi_{f}(y^{[k]}|y^{[1:k-1]},x)$. We denote the best response as $y^{*}$, that is, the ground truth response. Ideally, a model’s performance could be measured by its loss relative to this ground truth, for instance, $L_{f}(\hat{y})=\|y^{*}-\hat{y}\|$. However, as not all tasks have an accessible ground truth, a different metric is required. Therefore, we also use the uncertainty metric following Huang et al. (2025) in our framework. We define the uncertainty for a response $\hat{y}$ given its prompt $x$ and a model $f$ as its negative log-likelihood: $U_{f}(\hat{y})=-\log\pi_{f}(\hat{y}|x).$ The model’s capability is inversely related to the uncertainty: a lower uncertainty signifies a higher capability. Although low uncertainty is often correlated with high capability (i.e., accuracy), they are not conceptually equivalent.
In this framework, we use uncertainty as a proxy metric for model capability, a common practice in related literature (Huang et al., 2025). The validity of this proxy in our context is empirically supported by our experiments (e.g., Figure 3), which show that as the model’s uncertainty decreases during self-improvement, its task accuracy concurrently increases. Therefore, throughout this paper, when we refer to an increase in capability, it should be understood as a decrease in uncertainty.
@@
$\hat{y}_{i}^{\text{BoN}}=\mathop{\mathrm{argmin}}_{\{\hat{y}_{i,j}:s(\hat{y}_{i,j})\geq\sigma\}}\frac{1}{L(\hat{y}_{i,j})}U_{f}(\hat{y}_{i,j}|x_{i}),$ (4)
@@
Self-Improvement. We deploy self-improvement based on the above solver-verifier framework, similar to Huang et al. (2025). Notably, the verifier is slightly different, since we want to include more verification information. Overall, we first generate responses from the LLM. The optimization objective is to minimize the average uncertainty of BoN responses. The loss function $L_{t}(f)$ for a training step $t$ with function $f$ is defined as the verifier uncertainty $U_{v}(t)$:
$L_{t}(f)\triangleq U_{v}(t)=-\frac{1}{n}\sum_{i=1}^{n}\log\ \pi_{f}(\hat{y}_{i}^{\text{BoN}}(t)|x_{i}).$ (6)
By minimizing $L_{t}(f)$, we steer the model to increase the likelihood of generating high-quality responses, improving its solver capability. We summarize one-step self-improvement algorithm in Algorithm 1.
Algorithm 1 Self Improvement (One Step)
[…]
2:for each prompt $x_{i}\in\mathcal{X}$ do
3:  $C_{x_{i}}\leftarrow\emptyset$;
4:  for $j\leftarrow 1$ to NN do
5:   Generate response $\hat{y}_{i,j}\sim\pi_{f}(\cdot|x_{i})$;
6:   Ask model to evaluate $\hat{y}_{i,j}$ and return a score $s(\hat{y}_{i,j})$;
7:   if $s(\hat{y}_{i,j})\geq\sigma$ then
8:    Append $\hat{y}_{i,j}$ to $C_{x_{i}}$;
9:   end if
10:  end for
11:  $\hat{y}_{i}^{\text{BoN}}\leftarrow\mathop{\mathrm{argmin}}_{C_{x_{i}}}\frac{1}{L(\hat{y}_{i,j})}U_{f}(\hat{y}_{i,j}|x_{i})$;
12:  Append $\hat{y}_{i}^{\text{BoN}}$ to $\mathcal{Y}^{\text{BoN}}$;
13:  Append $U_{f}(\hat{y}_{i}^{\text{BoN}})$ to $\mathcal{U}^{\text{BoN}}$;
14:end for
[…]
16: Self-improved model $\hat{f}$
@@ 3.2 Self-Improvement Dynamics
We assume that the change of the solver and verifier capability is driven by a gap potential energy $E(t)$, expressed as a function of the Capability Gap $G(t)$ (Equation 7), namely $E(t)=f(G(t))$, where $f(G)$ is a differentiable and monotonically increasing function for $G\geq 0$, with its minimum $f(0)=0$. In this paper, we simply use uncertainty to evaluate the response quality, leading the capability gap defined with the uncertainty, as discussed in Section 3.1. We assume that the solver capability and the verifier capability both increase during the process of self-improvement, which is widely observed in the related works (Song et al., 2025). With the analysis above, following the concept of potential energy, the theoretical framework starts with the following assumptions:
@@
where $\alpha,\beta\geq 0$ are coefficients related to the decreasing rate of $U_{s}(t),U_{v}(t)$. In this framework, Equation 8 represents the initial conditions. Since for LLM, the verification capability usually outperforms the solver capability in real-world applications, we assume that $U_{s,0}>U_{v,0}$Equation 9 represents the iterative conditions, where we assume that $\alpha>\beta$, indicating that the solver capability increases faster than the verifier. Based on the above assumptions, we derive the differential equation governing the Capability Gap:
@@
However, for a non-linear function $f(G)$, a closed-form solution for above integral is not guaranteed. To simplify this problem, we approximate the potential energy function $E(t)$ with a linear form, $E(t)\approx kG(t)-b$, derived from a first-order Taylor expansion. Figure 2 illustrates the strong linear relationship between uncertainty gap $G$ and its rate of change $dG/dt=-(\alpha-\beta)E(t)$ on Phi-4-mini with QE method, which provides robust evidence for our approximation. Experiment results on other models are presented in Figure 7. Based on the Equation 9, we derive proposition 3.1:
@@
Equation 11 and Equation 12 demonstrate that uncertainties $U_{s}$, $U_{v}$, and their gap $G(t)$ all converge exponentially toward their respective limits. The proof of proposition 3.1 is shown in Section B.1. We derive two key corollaries from proposition 3.1.
Corollary 3.1.
@@
The proof of corollary 3.1 is shown in Section B.2corollary 3.1 shows that a larger initial verifier-solver gap ($G_{0}$) leads to a lower final solver uncertainty ($U_{s,\infty}$). This aligns with the core intuition that the performance gap is the driver of the self-improvement process.
Corollary 3.2.
@@
The proof of corollary 3.2 is shown in Section B.3corollary 3.2 provides theoretical guidance for selecting the total number of training epochs $T$ in practical application. In practice, we first estimate the key convergence parameters by training for a few initial epochs and fitting the early-stage performance data. These parameters can then be used to determine an appropriate total training time $T$ to balance performance with computational costs.

### v4

# Theoretical Modeling of Large Language Model Self-Improvement Training Dynamics Through Solver-Verifier Gap

@@ Abstract
Self-improvement is a significant techniques within the realm of large language model (LLM), aiming to enhance the LLM performance without relying on external data. Despite its significance, generally how LLM performance evolves during the self-improvement process remains underexplored. In this paper, we theoretically model the training dynamics of self-improvement via the concept of solver-verifier gap. This is inspired by the conjecture that the performance enhancement of self-improvement stems from the gap between LLM’s solver capability and verifier capability. Based on the theoretical framework, we further show how to model the entire training trajectory. This framework allows quantifying the capability limit of self-improvement by fitting the theoretical model to the experiment results. We validate the effectiveness of the theoretical framework on various LLMs and datasets. Beyond self-improvement, we extend our analysis to investigate how external data influences these dynamics within the framework. Notably, we find that under limited external data regimes, such external data can be utilized at any stage without significantly affecting final performances, which accords with the empirical observations.
@@ 1 Introduction
In this paper, we theoretically model the training dynamics of LLM self-improvement, inspired by the conjecture that self-improvement capability arises from the gap between an LLM’s solver capability and verifier capability  [^1] We term this the solver-verifier gap to ensure consistency with our mathematical notation ($U_{s}$ for Solver, $U_{v}$ for Verifier) and our specific Best-of-N implementation. We explicitly acknowledge that this concept builds directly upon the generation-verification gap originally conceptualized by Song et al. (2025), and we adopt this terminology with deep respect for their foundational contribution.. Specifically, we define these two capabilities as follows:
@@
where $\alpha^{\prime},\beta^{\prime},k$ denotes the constants with detailed formulations in Section˜3.2, and $U_{s,\infty},U_{v,\infty}$ represent the solver capability and the verifier capability at convergence, respectively. Along the trajectory, both capabilities follow exponential convergence laws according to the framework. Besides, we specifically focus on the solver’s ultimate capability $U_{s,\infty}=\frac{1}{\alpha-\beta}(\alpha U_{v,0}-\beta U_{s,0}+\alpha\frac{b}{k})$. This result reveals that the solver’s ultimate capability relates to both the initial capability $U_{s,0},U_{v,0}$ and the coefficient $\alpha,\beta$. Therefore, the ultimate capability might be improved given a larger verifier-solver gap at initialization, in accordance with our insights. Detailed derivations are provided in Section˜3.2.
@@
From the experimental perspective, we find in Section˜4 that such theoretical modeling demonstrates strong practical efficacy. Specifically, experimental results across various models and datasets in Figure˜3 reveal that the capability dynamics indeed follow an exponential law, as implied by the theoretical framework in Section˜3. Besides, we observe in Figure˜3 that the verifier capability consistently outperforms the solver throughout the self-improvement process, which might be empirical evidence of the solver-verifier gap’s crucial role in driving capability improvement. To further investigate the role of the solver-verifier gap, we conduct experiments in Section˜4.3 on multiple LLMs demonstrating that the gap generally occurs in practice. These findings hold across varying sample sizes and even under cross-evaluation scenarios.
@@
We further analyze in Section˜5 the application of this framework in cross-improvement, a potential approach to enhance the ultimate capability of self-improvement. Cross-improvement in this paper refers to the utilization of external data in the verification step (details in Figure˜7). Within the theoretical framework, we contend that cross-improvement outperforms self-improvement due to the enhanced verification capabilities. We then derive the training dynamics under the cross-improvement regimes, with enhanced verification capabilities compared to self-improvement. This analysis can further assist in answering the question: given the limited amount of external data, how should one allocate these external data during the cross-improvement training process? Our theoretical analyses demonstrate that the timing of using external data is not crucial; rather, the utilization of such data genuinely enhances the training process. Therefore, external data can be incorporated at any stage as desired. Experiments in Section˜5.2 on the cross-improvement validate the theoretical findings. Our main contributions are presented in Section˜A.2.
@@ 2 Related Work
Self-Improvement. Self-improvement aims to enhance model performance without relying on external information (Huang et al., 2023; Wang et al., 2023; Bai et al., 2022; Pang et al., 2024). Self-improvement is of significant importance as it enables models to adapt and evolve autonomously, thereby facilitating their effectiveness across a wide range of real-world scenarios, such as reasoning (Zelikman et al., 2022; Peng et al., 2024; Huang et al., 2023), alignment (Wang et al., 2023; 2024b; Ding et al., 2024), and planning (Tian et al., 2024). In this paper, we consider a branch of self-improvement that fine-tunes with the output of LLM (Amini et al., 2024; Sessa et al., 2024; Gui et al., 2024; Pace et al., 2024; Ouyang et al., 2022; Rafailov et al., 2023). Alternative approaches to self-improvement include self-distillation (Buciluǎ et al., 2006; Hinton et al., 2015; Zhang et al., 2019) which involves transferring knowledge from a larger, more complex model to a smaller one, and self-correction (Kumar et al., 2024; Liu et al., 2024) where the model identifies and rectifies its own errors, etc. A body of research has also explored the potential negative consequences of self-improvement, including degradation issues (Bertrand et al., 2024; Gerstgrasser et al., 2024) and failures on out-of-domain reasoning tasks (Yuan et al., 2025).
Theoretical Understandings on Self-Improvement. Theoretical insights into self-improvement could potentially enhance comprehension, thereby making self-improvement more reliable (Yampolskiy, 2015). Previous works have theoretically studied self-improvement via self-distillation techniques, providing convergence rates for linear models (Mobahi et al., 2020; Frei et al., 2022; Das & Sanghavi, 2023; Pareek et al., 2024), neural networks (Allen-Zhu & Li, 2023), and general models (Boix-Adsera, 2024). In the realm of LLM, several works have theoretically explored self-improvement with in-context alignment (Wang et al., 2024c), reinforcement learning (Talvitie, 2017; Choi et al., 2024; Gandhi et al., 2025), meta learning (Kirsch & Schmidhuber, 2022), and diffusion models (Fu et al., 2024). Nevertheless, the theoretical understanding of self-improvement in the context of LLM training dynamics remains underexplored.
@@ 3 Theoretical Modeling of Self-Improvement Training Dynamics
This section presents a theoretical framework for sketching training dynamics in LLM self-improvement. We start by introducing necessary notations of self-improvement in Section˜3.1. We then theoretically model the self-improvement dynamics in Section˜3.2.
@@ 3.1 Solver and Verifier
Notations. Let $(x,y)$ denote a prompt-response pair, with $y^{[k]}$ denoting the $k$-th token and $y^{[1:k]}$ denoting the first $k$ tokens. We use $L(y)$ to represent the length of $y$. Let $\pi_{f}(y|x)$ denote the probability that a model $f$ generates a response $y$ given a prompt $x$, where $\pi_{f}(y|x)$ can be split with the auto-regressive structure of the response $\pi_{f}(y|x)=\prod_{k=1}^{L(y)}\pi_{f}(y^{[k]}|y^{[1:k-1]},x)$. We denote the best response as $y^{*}$, that is, the ground truth response. Ideally, a model’s performance could be measured by its loss relative to this ground truth, for instance, $L_{f}(\hat{y})=\|y^{*}-\hat{y}\|$. However, as not all tasks have an accessible ground truth, a different metric is required. Therefore, we also use the uncertainty metric following Huang et al. (2025) in our framework. We define the uncertainty for a response $\hat{y}$ given its prompt $x$ and a model $f$ as its negative log-likelihood: $U_{f}(\hat{y})=-\log\pi_{f}(\hat{y}|x).$ The model’s capability is inversely related to the uncertainty: a lower uncertainty signifies a higher capability. Further discussion is available in Section˜A.3.
@@
$\hat{y}_{i}^{\text{BoN}}=\operatorname*{arg\,min}_{\{\hat{y}_{i,j}:s(\hat{y}_{i,j})\geq\sigma\}}\frac{1}{L(\hat{y}_{i,j})}U_{f}(\hat{y}_{i,j}|x_{i}),$ (4)
@@
Self-Improvement. We deploy self-improvement based on the above solver-verifier framework, similar to Huang et al. (2025). Notably, the verifier is slightly different, since we want to include more verification information. Overall, we first generate responses from the LLM. The optimization objective is to minimize the average uncertainty of BoN responses. The loss function $L_{t}(f)$ for a training step $t$ with function $f$ is defined as the verifier uncertainty $U_{v}(t)$: $L_{t}(f)\triangleq U_{v}(t)=-\frac{1}{n}\sum_{i=1}^{n}\log\ \pi_{f}(\hat{y}_{i}^{\text{BoN}}(t)|x_{i}).$ By minimizing $L_{t}(f)$, we steer the model to increase the likelihood of generating high-quality responses, improving its solver capability. We summarize one-step self-improvement algorithm in Algorithm˜1.
@@ 3.2 Self-Improvement Dynamics
We assume that the change of the solver and verifier capability is driven by a gap potential energy $E(t)$, expressed as a function of the Capability Gap $G(t)$ (Equation˜6), namely $E(t)=f(G(t))$, where $f(G)$ is a differentiable and monotonically increasing function for $G\geq 0$, with its minimum $f(0)=0$. In this paper, we simply use uncertainty to evaluate the response quality, leading the capability gap defined with the uncertainty, as discussed in Section˜3.1. We assume that the solver capability and the verifier capability both increase during the process of self-improvement, which is widely observed in the related works (Song et al., 2025). With the analysis above, following the concept of potential energy, the theoretical framework starts with the following assumptions:
@@
where $\alpha,\beta\geq 0$ are coefficients related to the decreasing rate of $U_{s}(t),U_{v}(t)$Equation˜7 represents the initial conditions. Since for LLM, the verification capability usually outperforms the solver capability in applications, we assume that $U_{s,0}>U_{v,0}$Equation˜8 represents the iterative conditions, where we assume that $\alpha>\beta$, indicating that the solver capability increases faster than the verifier. Based on the above assumptions, we derive the differential equation governing the Capability Gap:
@@
However, for a non-linear function $f(G)$, a closed-form solution for above integral is not guaranteed. To simplify this problem, we approximate the potential energy function $E(t)$ with a linear form, $E(t)\approx kG(t)-b$, derived from a first-order Taylor expansion. Figure˜2 illustrates the strong linear relationship between uncertainty gap $G$ and its rate of change $dG/dt=-(\alpha-\beta)E(t)$ on Phi-4-mini with QE method, which provides robust evidence for our approximation. Experiment results on other models are presented in Figure˜8. Based on the Equation˜8, we derive Proposition˜3.1:
@@
Equation˜10 and Equation˜11 demonstrate that uncertainties $U_{s}$, $U_{v}$, and their gap $G(t)$ all converge exponentially toward their respective limits. The proof of Proposition˜3.1 is shown in Section˜B.1. We derive two key corollaries from Proposition˜3.1.
Corollary 3.1
@@
The proof of Corollary˜3.1 is shown in Section˜B.2Corollary˜3.1 shows that a larger initial verifier-solver gap ($G_{0}$) leads to a lower final solver uncertainty ($U_{s,\infty}$). This aligns with the core intuition that the performance gap is the driver of the self-improvement process.
Corollary 3.2
@@
The proof of Corollary˜3.2 is shown in Section˜B.3Corollary˜3.2 provides theoretical guidance for selecting the total number of training epochs $T$ in practical application. In practice, we first estimate the key convergence parameters by training for a few initial epochs and fitting the early-stage performance data. These parameters can then be used to determine an appropriate total training time $T$ to balance performance with computational costs.